You are walking away from home at a constant velocity of 5 km per hour. The first landmark you see is your favorite fountain which happens to be 2 km from your house. Which of the following sequences describes the distances you cover at intervals of one hour, starting when you pass the fountain?
A. 5, 7, 9, 11, 13, ...
B. 2, 7, 12, 17, 22, ...
C. 2, 4, 6, 8, 10, ...
D. 10, 12, 14, 16, 18, ...
E. 5, 10, 15, 20, 25, ...

Answer by Theo(13342)

you're traveling at 5 kmph (kilometers per hour).

so every hour you will be 5 km further down the road.

you start at 2 km.

2 km + 5 = 7 + 5 = 12 + 5 = 17 + 5 = 22, etc.

your sequence would be selection B. 2, 7, 12, 17, 22, ...

there are the distances you would be away from home every hour, starting from 2 km away which is where the fountain is.

your 0 hour point is 2 km from home.
your 1 hour point is 7 km from home.
your 2 hour point is 12 km from home.
etc.